~Include a normal curve/t distribution curve.
~Your final answer must be a concluding statement.

To determine the views of students on a particular campus about palabok, an 11-point attitude scale was administered to a random sample of 40 students. This survey yielded a sample of 6 and a standard deviation of 1.5
1. Compute for the standard error of the mean.
2. Find the 95% confidence interval for the population mean.
3. Find the 99% confidence interval for the population mean.
1
Expert's answer
2020-10-26T16:59:04-0400

The provided sample mean is "\bar X = 6"

 and the sample standard deviation is s = 1.5. The size of the sample is n = 40 and the required confidence level is 95%.


1) standard error= "\frac{tc*s}{\sqrt{n}}"

Based on the provided information, the critical t-value for α=0.05 and df = n-1= 40-1=39

is tc​=2.023


so standard error= "\frac{2.023*1.5}{\sqrt{40}}" =0.48


ii)The 95% confidence for the population mean "\mu" is computed using the following expression

"[\overline{X} - \frac{tc*s}{\sqrt{n}},\overline{X} +\frac{tc*s}{\sqrt{n}}]"

Substitute the values ,

"[6 - \frac{2.023*1.5}{\sqrt{40}},6 + \frac{2.023*1.5}{\sqrt{40}}]"

=(6−0.48,6+0.48)

=(5.52,6.48)

iii)Based on the provided information, the critical t-value for α=0.01 and df = n-1= 40-1=39 is tc​=2.708

so confidence interval is given by

"[\overline{X} - \frac{tc*s}{\sqrt{n}},\overline{X} +\frac{tc*s}{\sqrt{n}}]"

"[6 - \frac{2.708*1.5}{\sqrt{40}},6 + \frac{2.708*1.5}{\sqrt{40}}]"

=(6−0.642,6+0.642)

=(5.358,6.642).